# thanks: These authors contribute equallythanks: These authors contribute equally

### v1

@@ Abstract
We subsequently assess multi-step-ahead forecasting, in which predictability is constrained by the rapid error growth inherent to the chaotic atmospheric system Kalnay (2003); Palmer (1993). Here, a direct multi-step forecast strategy is employed Taieb et al. (2012), wherein a separate model is trained for each forecast horizon $h$: ${\bf y}_{k+h}=f_{h}({\bf y}_{k},{\bf y}_{k-1},{\bf y}_{k-2},\ldots)$. This is implemented through multitasking in QRC, where distinct readout weights are trained for each target sequence $\{{\bf y}_{k+h}\}$ while utilizing the same input sequence $\{{\bf y}_{k}\}$. To evaluate the potential quantum advantages, we compare the performance of QRC with that of echo state networks (ESNs) Jaeger and Haas (2004), with results displayed in Fig. 4(b). ESNs with 500, 1000, 5000, and 10000 nodes (denoted as ESN($m$) for $m$-node networks) are evaluated, showing improved performance with increasing size until saturation, where the curves of ESN(5000) and ESN(10000) nearly overlap, indicating diminishing returns. For temperature forecasting, QRC attains accuracy comparable to that of ESN(1000), and even exceeds its average at larger forecast horizons $h$. Moreover, incorporating nonlinear post-processing may further enhances QRC’s performance: support vector regression (SVR) with a radial basis function (RBF) kernel Awad et al. (2015) enables QRC to achieve higher accuracy than ESN(10000), whereas ESNs gain no similar improvementsup . We attribute QRC’s strong performance to the superior expressibility Sim et al. (2019) and unique information processing capabilities Kubota et al. (2023) inherent to quantum models. Although humidity prediction is inherently more challenging due to its environmental sensitivity, QRC still achieves performance comparable to that of ESN(10000). These results show the capability of quantum systems to address complex dynamical prediction tasks, highlighting the potential quantum advantages in real-world applications with current NISQ devices.
@@
Conclusion.—Building on the natural dynamics of a spin system, we develop a hardware-efficient quantum reservoir computing scheme that enables accurate time-series predictions in practical machine learning tasks. Implemented on a 9-spin system with time-multiplexed readouts, our approach achieves one to two orders of magnitude lower error compared to previous QRC experiments on NARMA benchmarks. In long-term weather forecasting, our quantum reservoir achieves higher prediction accuracy than classical reservoirs with thousands of nodes, suggesting that practical quantum advantages in time-series prediction may be attainable with current quantum hardware. We anticipate the computational capacity of our QRC protocol can be further improved by increasing the complexity of the reservoir dynamics, for example through pulse-engineered refinement of internal interactions, or by scaling to larger quantum processors. Moreover, integrating QRC with many-body phenomena, including dynamical phase transitions Mart´ınez-Peña et al. (2021); Xia et al. (2022); Kobayashi and Motome (2023, 2025), quantum scar states Serbyn et al. (2021); Bravo et al. (2022), and time crystals Zhang et al. (2017), represents a promising direction for future research, for it could help uncover physical principles to improve QRC performance systematically.
Acknowledgements.
@@
This work is supported by Innovation Program for Quantum Science and Technology (Grants No. 2024ZD0302000, 2021ZD0303205, 2024ZD0300100), National Natural Science Foundation of China (Grants No. T2388102, 12261160569, 92165108), National Basic Research Program of China (Grants No. 2021YFA1400900), Shanghai Municipal Science and Technology (2019SHZDZX01, 24DP2600100), and the XPLORER Prize. The authors acknowledge helpful discussion with the Anhui Meteorological Bureau of China.
@@
In this section we present the experimental setup employed in this work. Our experiments were conducted using a Bruker Avance III 400 MHz spectrometer at 298 K. The quantum reservoir is a 9-spin coupling system consisting of four carbon nuclei and five proton nuclei in a 13C-labeled crotonic acid molecule, the structure of which is shown in TABLE 2. The methyl protons $\rm H_{3}$, $\rm H_{4}$ and $\rm H_{5}$ are chemically and magnetically equivalent and therefore indistinguishable. TABLE 2 lists the chemical shifts $\nu_{i}$ (diagonal elements) and the $J$-coupling constants $J_{ij}$ (off-diagonal elements), along with the corresponding relaxation times. In particular, the $T_{2}$ times account for dephasing induced by magnetic field inhomogeneities (referred to as $T_{2}^{*}$ in NMR Levitt (2008)).
@@
Table 2: Sample parameters and molecular structure of nine-spin crotonic acid sample. The left portion presents chemical shifts (diagonal elements) and $J$-couplings (off-diagonal elements), while the right lists the corresponding relaxation times.
@@
The preprocessed input component is encoded in the pulse area of a radiofrequency (RF) field, with its phase aligned along the $x$-axis and frequency resonant with designated nuclei (in this work we consider inputs with ${\rm dim}\leq 2$, while higher‑dimensional inputs can be encoded using general parameterized circuits Li et al. (2022); Caro et al. (2021)). Each pulse, with a duration of approximately $10^{-5}$ s, is short enough that the system’s free evolution during the pulse can be neglected, thereby effectively implementing a global rotation on all proton spins. Although slight deviations in the rotation angle and axis may arise from detuning between the chemical shift and the RF frequency, this effect remains fixed and is incorporated within the quantum reservoir. The pulses are applied at regular intervals $\tau$, during which the system evolves under its natural Hamiltonian and relaxation dynamics without external control. The readout is performed at the end of the evolution period to measure the protons, since the signal-to-noise ratio for carbon is significantly lower. Consequently, the carbon spins are treated as an inaccessible bath, demonstrating that the computational power of the system can be harnessed solely by accessing a subset of its spins. Two types of readout schemes are implemented: one scheme involves measuring the observable expectations $\langle\sigma^{m}_{x,y,z}\rangle$ with $m={\rm H_{1}},{\rm H_{2}}$, and the methyl protons, while the time multiplexing readout scheme is described in the subsequent section. The readout vector dataset is normalized using Z-score standardization followed by ridge regression, with the regularization strength determined via 10-fold cross validation.
@@
In our experiments, the number of washout steps $n_{\rm wo}$ was determined by extending the washout sequence until the NMR spectral readout (see FIG. 5(b)) remained invariant across different initial states. For NARMA tasks, we used $n_{\rm wo}=100$ for $\tau=$0.3\text{\,}\mathrm{s}$\$ and 1000 for $\tau=$0.01\text{\,}\mathrm{s}$\$, which may be larger than the minimum required to achieve convergence. As described in the main text, a rewinding protocol was employed in our experiments Mujal et al. (2023); that is, each experiment is reinitialized with last $n_{\rm wo}$ input-evolution cycles (from $k-n_{\rm wo}$ to k1k-1) to reprepare the reservoir to $\rho_{k-1}$, then perform the input-evolution-readout procedure for the $k$-th step. Other experimental protocols have been reported in recent years Mujal et al. (2023); Kobayashi et al. (2024), such as the weak measurement protocol that enables online data processing without restarting experiments, which we plan to explore in future studies.
.2 II.    FID signal and NMR spectrum as readout
@@
Here we outline the mathematical details. Assume that a unitary transformation $U$ is applied to the final state $\rho$ to be measured, and the FID signal is subsequently recorded from the first qubit. This signal can be expressed as (with $\omega_{i}=2\pi\nu_{i}$):
@@
Figure 5: (a) The real part of the free induction decay (FID) signal obtained from the thermally equilibrated proton spins of crotonic acid, truncated to show the first 4096 data points (half of the total acquisition time), with the signal amplitude enhanced by the spectrometer’s receiver amplification. (b) The real part of corresponding NMR spectrum, produced by Fourier transforming the FID signal. The intense peak near 1880 Hz originates from $\rm H_{2}O$ in the solvent. The red, green, and orange boxes highlight the spectral regions of crotonic acid’s $\rm H_{1}$, $\rm H_{2}$, and methyl protons $\rm H_{Me}$ (i.e. the equivalent $\rm H_{3},H_{4},H_{5}$), respectively; the insets display magnified views of these regions. Although the signal intensities of the crotonic acid protons are two to three orders of magnitude lower than that of the water peak, they still exhibit sufficiently high signal-to-noise ratios.
Thus far, we have explicitly detailed the mechanism for extracting state information via time-multiplexing (i.e., the FID signal) under a liquid NMR Hamiltonian, while the more general process is to learn an implicit linear transformation between the temporal measurements and the state representation Yang et al. (2020); Chantasri et al. (2020); Xiao et al. (2024). However, in practice, we use the spectral peaks of the FID signal after Fourier transformation as readout features, as highlighted in FIG. 5(b). This choice is motivated by the fact that the FID signal contains a significant contribution from $\rm H_{2}O$ in the solvent, which produces a dominant peak near 1880 Hz. Moreover, removing signal-independent components (e.g. random noise) can reduce computational complexity and mitigate the risk of overfitting. The spectral peaks can be described analytically as:
$S(\omega)=\int_{0}^{\infty}S(t)e^{i\omega t}{\rm d}t\propto\sum_{\bf q}M_{\bf q}\left[\frac{\lambda}{\left(\omega-\Omega_{\bf q}\right)^{2}+\lambda^{2}}+i\frac{\omega-\Omega_{\bf q}}{\left(\omega-\Omega_{\bf q}\right)^{2}+\lambda^{2}}\right]=\sum_{\bf q}M_{\bf q}\left[A_{\bf q}(\omega)+iD_{\bf q}(\omega)\right],$ (7)
where $A_{\mathbf{q}}(\omega)$ and $D_{\mathbf{q}}(\omega)$ correspond to the absorption and dispersion Lorentzian line shapes, respectively, centered at $\Omega_{\bf q}$ with a linewidth of $\lambda$. Typically, one would extract each component by fitting the spectrum; however, for samples with many spins, overlapping peaks (i.e. when $|\Omega_{\bf q}-\Omega_{\bf q^{\prime}}|\lesssim\lambda$) complicate such fitting. Only single-qubit observable expectations can be obtained by integrating over the corresponding spectral region. Our method, by contrast, directly utilizes the NMR spectral peaks as readouts without explicit post-processing, instead automatically extracting the necessary state information via linear regression. The effectiveness of this approach has been validated through both simulations (Section .5) and experiments (main text). Notably, no classical nonlinear enhancement techniques were introduced in this process. In our implementation, the truncated spectral peaks are merged into 653-dimensional features. While further dimensionality reduction is possible, given the observation that adjacent points on a Lorentzian peak exhibit high linear dependence, such reduction would entail a trade-off between performance and simplicity that we have not explored in detail.
.3 III.    Numerical characterization of spin-based quantum reservoir dynamics
In this section, we presents the numerical implementation of the input injection and the subsequent system evolution in the quantum reservoir. The normalized input $\bar{s}_{k}\in[0,1]$ is applied via a pulse on a specific nuclear channel (e.g., proton), which is equivalent to performing global rotations on the corresponding spins:
$U_{s_{k}}=\prod_{n\in S_{\rm in}}R^{(n)}_{x}\bigl{(}\arcsin(\bar{s}_{k})\bigr{)},$ (8)
where $S_{\rm in}$ denotes the set of qubits constituting the input channel.
The evolution of the system is governed by the Lindblad master equation:
[…]
with the Hamiltonian
We next characterize the relaxation channel $\mathcal{R}$ in detail.
For a single spin, the relaxation channel is given by
with the collapse operators defined as
$\rho_{\rm eq}=\frac{1}{2}\begin{bmatrix}1+p&0\\ 0&1-p\end{bmatrix}$, (13)
at a rate $T_{1}^{-1}$, where $p$ represents the spin polarization (with $p=1$ recovering the standard amplitude damping channel) Nielsen and Chuang (2001). The operator $L_{z}$ accounts for pure dephasing due to spin-spin interactions and field inhomogeneities, with a characteristic time $T_{\phi}$. Combined with amplitude damping, the effective decoherence rate is given by
$T_{2}^{-1}=\left(2T_{1}\right)^{-1}+T_{\phi}^{-1}.$ (14)
$\mathcal{R}[\rho]=\sum_{n=1}^{N}\mathcal{R}^{(n)}[\rho],$ (15)
.4 IV.    The NARMA task setup and experimental prediction results

### v2

@@ Abstract
We subsequently assess multi-step-ahead forecasting, in which predictability is constrained by the rapid error growth inherent to the chaotic atmospheric system [31, 62]. Here, a direct multistep forecast strategy is employed [78], wherein a separate model is trained for each forecast horizon $h$: ${\bf y}_{k+h}=f_{h}({\bf y}_{k},{\bf y}_{k-1},{\bf y}_{k-2},\ldots)$. This is implemented through multitasking in QRC, where distinct readout weights are trained for each target sequence $\{{\bf y}_{k+h}\}$ while utilizing the same input sequence $\{{\bf y}_{k}\}$. To evaluate the potential quantum advantages, we compare the performance of QRC with that of echo state networks (ESNs) [28], with results displayed in Fig. 4(b). ESNs with 500, 1000, 5000, and 10000 nodes (denoted as ESN($m$) for $m$-node networks) are evaluated, showing improved performance with increasing size until saturation, where the curves of ESN(5000) and ESN(10000) nearly overlap, indicating diminishing returns. For temperature forecasting, QRC attains accuracy comparable to that of ESN(1000), and even exceeds its average at larger forecast horizons $h$. Moreover, incorporating nonlinear postprocessing may further enhance QRC’s performance: support vector regression (SVR) with a radial basis function (RBF) kernel [4] enables QRC to achieve higher accuracy than ESN(10000), whereas ESNs gain no significant improvement [1]. We attribute QRC’s strong performance to the superior expressibility [75] and unique information processing capabilities [41] inherent to quantum models. As humidity is strongly influenced by rapidly varying local processes such as convection and precipitation, its prediction becomes inherently more challenging. Nonetheless, QRC still achieves performance comparable to that of ESN(10000). These results show the capability of quantum systems to address complex dynamical prediction tasks, highlighting the potential quantum advantages in real-world applications with current NISQ devices.
@@
Conclusion—Building on the natural dynamics of a spin system, we develop a hardware-efficient quantum reservoir computing scheme that enables accurate time-series predictions in practical machine learning tasks. Implemented on a 9-spin system with time-multiplexed readouts, our approach achieves 1 to 2 orders of magnitude lower error compared to previous QRC experiments on NARMA benchmarks. In long-term weather forecasting, our quantum reservoir achieves higher prediction accuracy than classical reservoirs with thousands of nodes, suggesting that practical quantum advantages in time-series prediction may be attainable with current quantum hardware. We anticipate that the computational capacity of our QRC protocol can be further improved by increasing the complexity of the reservoir dynamics, for example, through pulse-engineered refinement of internal interactions, or by scaling to larger quantum processors. Moreover, integrating QRC with many-body phenomena, including dynamical phase transitions [51, 84, 35, 36], quantum scar states [74, 8], and time crystals [87], represents a promising direction for future research, as it could help uncover physical principles to improve QRC performance systematically.
@@
Acknowledgments—This work is supported by Quantum Science and Technology-National Science and Technology Major Project (Grants No. 2024ZD0302000, No. 2021ZD0303205 and No. 2024ZD0300100), National Natural Science Foundation of China (Grants No. T2388102, No. 12261160569 and No. 92165108), National Basic Research Program of China (Grants No. 2021YFA1400900), Shanghai Municipal Science and Technology (No. 25TQ003, No. 2019SHZDZX01 and No. 24DP2600100), and the XPLORER Prize. The authors acknowledge helpful discussions with the Anhui Meteorological Bureau of China.
Data availability—The data and code that support the findings of this study are openly available [25].
@@ Appendix A Experimental setup
In this section we present the experimental setup employed in this work. Our experiments were conducted using a Bruker Avance III 400 MHz spectrometer at 298 K. The quantum reservoir is a 9-spin coupling system consisting of four carbon nuclei and five proton nuclei in a 13C-labeled crotonic acid molecule, the structure of which is shown in TABLE 2. The methyl protons $\rm H_{3}$, $\rm H_{4}$, and $\rm H_{5}$ are chemically and magnetically equivalent and therefore indistinguishable. TABLE 2 lists the chemical shifts $\nu_{i}$ (diagonal elements) and the $J$-coupling constants $J_{ij}$ (off-diagonal elements), along with the corresponding relaxation times. In particular, the $T_{2}$ times account for dephasing induced by magnetic field inhomogeneities (referred to as $T_{2}^{*}$ in NMR [45]).
@@
Table 2: Sample parameters and molecular structure of the nine-spin crotonic acid sample. The left portion presents chemical shifts (diagonal elements) and $J$-couplings (off-diagonal elements), while the right lists the corresponding relaxation times. The chemical shifts are given in the rotating frame, referenced to 100.6273 MHz for 13C and 400.2118 MHz for 1H.
@@
The preprocessed input component is encoded in the pulse area of a radiofrequency (RF) field, with its phase aligned along the $x$-axis and frequency resonant with designated nuclei (in this work we consider inputs with ${\rm dim}\leq 2$, while higher‑dimensional inputs can be encoded using general parameterized circuits). Each pulse, with a duration of approximately $10^{-5}$ s, is short enough that the system’s free evolution during the pulse can be neglected, thereby effectively implementing a global rotation on all proton spins. Although slight deviations in the rotation angle and axis may arise from detuning between the chemical shift and the RF frequency, this effect remains fixed and is incorporated within the quantum reservoir. The pulses are applied at regular intervals $\tau$, during which the system evolves under its natural Hamiltonian and relaxation dynamics without external control. The readout is performed at the end of the evolution period to measure the protons, since the signal-to-noise ratio for carbon is significantly lower. Consequently, the carbon spins are treated as an inaccessible bath, demonstrating that the computational power of the system can be harnessed solely by accessing a subset of its spins. Two types of readout schemes are implemented: one scheme involves measuring the observable expectations $\langle\sigma^{m}_{x,y,z}\rangle$ with $m={\rm H_{1}},{\rm H_{2}}$, and the methyl protons, while the time multiplexing readout scheme is described in the subsequent section. The readout vector dataset is normalized using Z-score standardization followed by ridge regression, with the regularization strength determined via 10-fold cross-validation.
@@
In our experiments, the number of washout steps $n_{\rm wo}$ was determined by extending the washout sequence until the NMR spectral readout (see FIG. 5(b)) remained invariant across different initial states. For NARMA tasks, we used $n_{\rm wo}=100$ for $\tau=$0.3\text{\,}\mathrm{s}$\$ and 1000 for $\tau=$0.01\text{\,}\mathrm{s}$\$, which may be larger than the minimum required to achieve convergence. As described in the main text, a rewinding protocol was employed in our experiments [57]; that is, each experiment is reinitialized with last $n_{\rm wo}$ input-evolution cycles (from $k-n_{\rm wo}$ to k1k-1) to reprepare the reservoir to $\rho_{k-1}$, then perform the input-evolution-readout procedure for the $k$-th step. Other experimental protocols have been reported in recent years [57, 34], such as the weak measurement protocol that enables online data processing without restarting experiments, which we plan to explore in future studies.
Appendix B Theoretical basis of time-multiplexed readout
Time-multiplexed readout provides an efficient way to extract information from a quantum system, particularly when direct access to many observables is not available. When performed during input-independent evolution, this process is conceptually analogous to dynamical quantum tomography, where information about the (input) state is inferred from a time series of measurements of fixed observables [33, 85]. By accessing Heisenberg-evolved observables at multiple evolution times, the time-multiplexed scheme effectively increases the number of independent readout variables, thereby raising the upper bound on the achievable performance.
Formally, consider a $d$-dimensional quantum system initialized in state $\rho$, whose dynamics are described by a quantum channel $\Phi_{t}(\rho)=\sum_{a}K_{t,a}\rho K_{t,a}^{\dagger}$. The expectation value of an observable $O$ at time $t$ is
$\langle O\rangle_{t}={\rm Tr}[\rho(t)O]={\rm Tr}\,\Big[\sum_{i}K_{t,a}\rho K_{t,a}^{\dagger}O\Big]={\rm Tr}\,\Big[\rho\sum_{i}K_{t,a}^{\dagger}OK_{t,a}\Big]={\rm Tr}[\rho\,O(t)],$ (6)
where $O(t)=\Phi_{t}^{\dagger}(O)$ denotes its Heisenberg-picture evolution. The accessible measurement space is thus $\mathcal{M}={\rm span}\{O(t)\,|\,O\in\mathcal{S},\,t\in\mathcal{T}\}$, with $\mathcal{S}$ the set of directly measurable independent observables and $\mathcal{T}$ the sampling times. We use ${\rm dim}(\mathcal{M})$ to denote the dimension of the operator space $\mathcal{M}$, and $|\mathcal{S}|$ to denote the number of observables in the set $\mathcal{S}$. When the dynamics generate nontrivial operator evolution with $K_{t,a},O]\neq 0$, the condition ${\rm dim}(\mathcal{M})>|\mathcal{S}|$ typically holds, indicating that time multiplexing provides access to a larger set of independent operator components than direct measurement alone. Each time-evolved operator $O(t)_{i}$ (where $i$ indexes distinct observable–time pairs) can be expanded in an operator basis as $O(t)_{i}=\sum_{k=1}^{{\rm dim}(\mathcal{M})}\alpha_{ik}E_{k}$, yielding the QRC output
$y=\sum_{i}w_{i}{\rm Tr}[\rho O(t)_{i}]+b=\sum_{k=1}^{{\rm dim}(\mathcal{M})}w_{k}^{\prime}\,{\rm Tr}[\rho E_{k}]+b,$ (7)
which allows the QRC to exploit these additional degrees of freedom.
In the ideal case with sufficient sampling points, ${\rm dim}(\mathcal{M})$ can reach its maximal value, given by the dimension of the Krylov subspace spanned by the Heisenberg-evolved observables $\{\Phi_{t}^{\dagger}(\tilde{O}_{i})\,|\,t\geq 0\}$. The Krylov dimension therefore sets an upper bound on the number of independent readout variables accessible via time multiplexing, and thus limits the reservoir’s total information processing capacity (IPC), which quantifies its ability to implement linear memory and nonlinear mappings of time-lagged inputs [16]. By increasing the number of independent measurements from $|\mathcal{S}|$ to ${\rm dim}(\mathcal{M})$, the time-multiplexed scheme effectively raises the upper bound on the total IPC and enhances the achievable performance of QRC.
When the generators of dynamics are known, ${\rm dim}(\mathcal{M})$ can be inferred analytically. Here we take the NMR setting used in this work as an example. Under the internal Hamiltonian $H=\sum_{i=1}^{N}\pi\nu_{i}\,\sigma_{z}^{i}+\sum_{i<j}\frac{\pi}{2}J_{ij}\sigma_{z}^{i}\sigma_{z}^{j}$, the transverse operator on the first spin $\sigma_{y}^{1}+i\sigma_{x}^{1}$ evolves in the Heisenberg picture as
@@ Appendix B Theoretical basis of time-multiplexed readout
which contains $2^{N-1}$ coherences with distinct oscillation frequencies and can be expanded into $2^{N}$ Pauli operators corresponding to multi-spin correlations. Such information is directly captured in the free-induction-decay (FID) signal, which we introduce in Section C. More generally, even without exact knowledge of the dynamics, it has been proven that under sufficiently generic (symmetry-free) evolution, a single observable sampled over time can span the full operator space [70], ensuring the effectiveness of the time-multiplexed scheme.
We finally analyze the experimental runtime of time-multiplexed readout in comparison with conventional single-time readout, i.e., measuring multiple specified observables of a static state.
  • (a)
    On platforms supporting continuous weak measurement, all sampling points $\{O(t_{i})\}$ can be obtained within a single experimental run, as the system–detector coupling is sufficiently weak to suppress measurement back-action. In this way, time-multiplexed readout extracts far more effective observables than a single-time measurement of $O$, enabling markedly improved performance without adding extra runtime. Such continuous monitoring schemes are available on various platforms [76, 59, 13, 14], including NMR systems, where the standard FID detection naturally provides a weak-measurement time trace.
  • (b)
    On platforms limited to projective measurements, obtaining $M$ sampling points requires $M$ separate experiments. For a fair comparison, single-time readout would also need approximately $M$ experiments to access a comparable number of independent observables, since QRC performance depends strongly on readout dimension [16]. Thus, the total experimental runtime of the two approaches is expected to be similar when matched to the same performance target. Parallel measurement of commuting observables does not change this conclusion, because the same strategy can be applied to each experiment in time-multiplexed readout. Moreover, time-multiplexed readout is often experimentally favorable, as it uses free evolution and avoids the need to implement dedicated control sequences for different observables.
Appendix C Experimental implementation of time-multiplexed readout using FID signal
@@
Figure 5: Feature extraction from the NMR free-induction-decay (FID) signal. (a) The real part of the FID signal obtained from the proton spins of crotonic acid, showing the first 2048 points (one quarter of the total acquisition). The signal amplitude is enhanced by the spectrometer’s receiver gain. (b) The real part of the corresponding NMR frequency spectrum, produced by Fourier transforming the FID signal. The intense peak near 1880 Hz originates from $\rm H_{2}O$ in the solvent. The red, green, and orange boxes highlight the spectral regions of crotonic acid’s $\rm H_{1}$, $\rm H_{2}$, and methyl protons $\rm H_{Me}$ (i.e., the equivalent $\rm H_{3},H_{4},H_{5}$), with insets showing magnified views. Although these signal intensities are two to three orders of magnitude lower than that of the $\rm H_{2}O$ peak, they still exhibit sufficiently high signal-to-noise ratios. (c) Normalized readout vector constructed by concatenating the spectral amplitudes from the highlighted regions, serving as time-multiplexed readout features for QRC.
In practice, we do not use the raw FID signal directly, but instead perform a Fourier transform and select the sparsely distributed spectral peaks as readout features, as shown in Fig. 5(b). This choice is motivated by two considerations. First, the time-domain FID contains a large background contribution from $\mathrm{H_{2}O}$ molecules in the solvent, which can be removed by spectral filtering in the frequency domain, since it appears as a well-isolated peak near 1880 Hz. Second, discarding frequency components that do not carry reservoir-related signals reduces the effective feature dimension, thereby lowering computational overhead and mitigating potential overfitting. For the selection of specific features, we take the real spectral amplitudes around the signal peak positions and concatenate them into a 653-dimensional readout vector, as shown in Fig. 5(c). This includes all major peaks while keeping the feature number low for stable regression. Small variations in the selected spectral range have a negligible influence on QRC performance. Further dimensionality reduction is possible because adjacent points within each spectral peak are highly correlated. However, this involves a trade-off between simplicity and predictive accuracy, which can be task-dependent and requires further analysis in future work.
Appendix D The NARMA task setup and experimental prediction results